# Non-Overlapping Placement of Macro Cells based on Reinforcement Learning in Chip Design

### v1

@@ Abstract
Index Terms:
Placement, macro cells, reinforcement learning, chip design
I Introduction
@@ III-B Action Space
III-C Rewards
In the placement process, an effective reward design is crucial because it directly affects the learning efficiency of the RL agent and the quality of the final layout. The reward design proposed in this paper carefully considers the critical performance indicators of chip layout, namely HPWL, layout density, and routing congestion (see Section IV-B), to ensure that the resulting layout meets physical requirements and electrical performance. The basic idea of the reward design is to guide the policy network in making the best macro cell placement decisions by precisely calculating the impact of each action on layout quality. The reward of SRLPlacer consists of two parts: one is the overall reward $R_{t}$ after the placement of all macro and standard cells has been completed, and the other is the immediate reward $R_{i}$ after one macro cell is placed. Together, these two rewards constitute the feedback for learning the policy network, aiming to achieve an optimal layout.
@@ III-C2 Immediate Reward
where $\alpha_{i}$ and $\beta_{i}$ are hyperparameters to balance the importance of immediate reward wire length and layout density; $WL$ and $H$ are as described in Section III-C1; $\tilde{P}$ represents the state of placed macro cells; $D$ is used to approximate the layout density for a given placement state.
@@ III-E Environment
IV Experiments
IV-A Benchmark and Settings
TABLE I: Statistics of ISPD-2005 benchmark dataset.
@@ IV-A Benchmark and Settings
TABLE II: The results of the ablation study (in HPWL ×$\times 10^{6}$).
@@
TABLE III: The results of the quantitative analysis (in HPWL ×$\times 10^{6}$).
@@
The validation experiments for SRLPlacer were conducted on the ISPD2005 benchmark[28]. This dataset was first introduced at the 2005 International Symposium on Physical Design (ISPD) and is mainly used for evaluating and comparing the performance of routing methods. ISPD2005 consists of different test cases, each representing a routing problem that must be solved. These test cases simulate the challenges in the chip design, including numerous netlists and complex routing areas, requiring the evaluated methods to effectively handle large-scale routing problems while optimizing critical paths to reduce total wire length and congestion. Although the ISPD2005 dataset was initially designed for routing tasks, its inclusion of high-density routing areas and complex constraints makes it a valuable resource for evaluating placement algorithms. Using the ISPD2005 dataset for placement experiments allows focusing on exploring how to effectively place macro and standard cells and how to deal with the impact of routing constraints on placement strategies. The statistics of the right test cases in ISPD2005 are shown in Table I.
[…]
IV-B Metrics
IV-C Ablation Study
[…]
IV-D Quantitative Analysis
In exploring the application of RL methods in macros placement tasks, SRLPlacer has demonstrated significant performance advantages, especially in handling macro cell placement problems. Compared with other popular placement methods, SRLPlacer successfully generated effective placement strategies and achieved the best performance in the critical performance metrics of HPWL across eight different scale and complexity test cases, as shown in Table III. This achievement can be attributed to SRLPlacer’s deep understanding and innovative solutions to both macro cell and standard cell placement problems, especially its ability to effectively avoid macro cell overlaps and optimize congestion and density, a significant challenge in traditional placement methods.
TABLE IV: Comparison of efficiency of different placement methods.
@@ IV-D Quantitative Analysis
The performance of SRLPlacer is reflected in its excellent result of HPWL and the significant improvement in placement efficiency, as shown in Table IV. This indicates that SRLPlacer is not only suitable for current chip design needs but also has the potential to meet future VLSI design challenges. Nevertheless, the application and development of SRLPlacer still face a series of challenges, including enhancing its adaptability to new chip designs and integrating more advanced optimization techniques. Future research directions could consider these challenges to promote the continuous optimization of SRLPlacer.
IV-E Visualized Layout

### v2

@@ Abstract
1 Introduction
@@ 3.2 Action Space
3.3 Rewards
In the placement process, an effective reward design is crucial because it directly affects the learning efficiency of the RL agent and the quality of the final layout. The reward design proposed in this paper carefully considers the critical performance indicators of chip layout, namely HPWL, layout density, and routing congestion (see Section 4.2), to ensure that the resulting layout meets physical requirements and electrical performance. The basic idea of the reward design is to guide the policy network in making the best macro cell placement decisions by precisely calculating the impact of each action on layout quality. The reward of SRLPlacer consists of two parts: one is the overall reward $R_{t}$ after the placement of all macro and standard cells has been completed, and the other is the immediate reward $R_{i}$ after one macro cell is placed. Together, these two rewards constitute the feedback for learning the policy network, aiming to achieve an optimal layout.
@@ 3.3.2 Immediate Reward
where $\alpha_{i}$ and $\beta_{i}$ are hyperparameters to balance the importance of immediate reward wire length and layout density; $WL$ and $H$ are as described in Section 3.3.1; $\tilde{P}$ represents the state of placed macro cells; $D$ is used to approximate the layout density for a given placement state.
@@ 3.5 Environment
4 Experiments
4.1 Benchmark and Settings
The validation experiments for SRLPlacer were conducted on the ISPD2005 benchmark[28]. This dataset was first introduced at the 2005 International Symposium on Physical Design (ISPD) and is mainly used for evaluating and comparing the performance of routing methods. ISPD2005 consists of different test cases, each representing a routing problem that must be solved. These test cases simulate the challenges in the chip design, including numerous netlists and complex routing areas, requiring the evaluated methods to effectively handle large-scale routing problems while optimizing critical paths to reduce total wire length and congestion. Although the ISPD2005 dataset was initially designed for routing tasks, its inclusion of high-density routing areas and complex constraints makes it a valuable resource for evaluating placement algorithms. Using the ISPD2005 dataset for placement experiments allows focusing on exploring how to effectively place macro and standard cells and how to deal with the impact of routing constraints on placement strategies. The statistics of the right test cases in ISPD2005 are shown in Table 1.
[…]
Table 1: Statistics of ISPD-2005 benchmark dataset.
@@ 4.1 Benchmark and Settings
Table 2: The results of the ablation study (in HPWL ×$\times 10^{6}$).
@@
Table 3: The results of the quantitative analysis (in HPWL ×$\times 10^{6}$).
@@
Table 4: Comparison of efficiency of different placement methods.
@@
4.2 Metrics
4.3 Ablation Study
[…]
4.4 Quantitative Analysis
In exploring the application of RL methods in macros placement tasks, SRLPlacer has demonstrated significant performance advantages, especially in handling macro cell placement problems. Compared with other popular placement methods, SRLPlacer successfully generated effective placement strategies and achieved the best performance in the critical performance metrics of HPWL across eight different scale and complexity test cases, as shown in Table 3. This achievement can be attributed to SRLPlacer’s deep understanding and innovative solutions to both macro cell and standard cell placement problems, especially its ability to effectively avoid macro cell overlaps and optimize congestion and density, a significant challenge in traditional placement methods.
The performance of SRLPlacer is reflected in its excellent result of HPWL and the significant improvement in placement efficiency, as shown in Table 4. This indicates that SRLPlacer is not only suitable for current chip design needs but also has the potential to meet future VLSI design challenges. Nevertheless, the application and development of SRLPlacer still face a series of challenges, including enhancing its adaptability to new chip designs and integrating more advanced optimization techniques. Future research directions could consider these challenges to promote the continuous optimization of SRLPlacer.
4.5 Visualized Layout